OpenStudy (smarty):

The Valley High School student council is sponsoring a dance. They anticipate that 500 students will attend. The ticket price is $3.00, and they estimate that for every $0.20 increase in ticket price, 25 fewer students will attend. What ticket price will maximize the student council’s profit?

OpenStudy (anonymous):

Let x be the number of groups of 25 students. Profit = price * quantity = (3 +0.2x) * (500 - 25x) = 1500 - 75x + 100x - 5x^2 = 1500 +25x - 5x^2 To find maximum profit, differentiate it. d(profit)/dx = -10x +25 Differentiate it once more to get -10, this shows that it is indeed a maximum point since it is less than zero. Let -10x+25 = 0 25= 10x x= 2.5 When x = 2.5, it is the maximum point. When x = 2.5 the ticket prize is, 3 +0.2(2.5) =3 + 0.5 = 3.5 Ticket prize is $3.50

OpenStudy (smarty):

Thanks, alot.!!!!

OpenStudy (anonymous):

Np. Hope you understand the workings, if you are unsure about anything, ask away. i will answer.

OpenStudy (smarty):

I got it! Work was very clear